[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Mon Oct 23, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Mon Oct 23 16:46:27 UTC 2017


TAFB Shift Log for Mon Oct 23 2017 12:46:24 GMT-0400 (EDT)

Shift: Day Shift

Pacific Forecaster: J Lewitsky

Atlantic Forecaster: C McElroy

Surface Analyst: A Latto

Technical Issues Summary:

GFE/NCP issues: 12z Pacific NWPS only came in for initial timestep despite  
run completing out to 168 hours. Mello investigating and checking to see if  
this is related to the security patch this morning.

N-AWIPS issues:

Other issues: Building AC System turned on with Fans running at highest  
level in all ops areas but warm/hot air blowing continuously. JL notified  
William at 1215pm. William was unaware of the issue but said that he would  
look into it. So far the server room is OK with 1 cooling unit at 69F,  
another at 70F, and a third slightly warm at 74F.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171023/641249a4/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list